OpenStudy (anonymous):

How do I find the Laplace Transform of e^(6t)?

OpenStudy (anonymous):

laplace of e ^at = 1/s-a so the answer is 1/s-6

OpenStudy (anonymous):

\[ \int_0^\infty e^{6t}dt \] Substitute u=6t and integrate. Then decide on how you have to pick the parameter to get a result that converges.